OpenStudy (anonymous):

The function h= is used to model fireworks being launched (or projected) into the air. The function models the height h after x seconds. a) How long will it take the fireworks to reach their greatest height? b) How high will the fireworks be their greatest height?

OpenStudy (anonymous):

\[h=-16x ^{2}+160x+4\]

OpenStudy (xapproachesinfinity):

to find the max height you need to take derivative

OpenStudy (trojanpoem):

Find maxima using derivative as xapproaches said then you will get the time put it in the main function and you will get the height.

OpenStudy (xapproachesinfinity):

\(\frac{dh}{dx}=-32x+160\) set this equal 0 to find the time it takes to reach maximum

OpenStudy (anonymous):

hello mam, i would like to know whether u are used to taking derivatives..?

OpenStudy (xapproachesinfinity):

you can actually discard calculus and do it with basic tools

OpenStudy (xapproachesinfinity):

that's just a parabola! and you learned that the prabolas has max at \(\left (\frac{-b}{2a}, h(\frac{-b}{2a})\right )\)

OpenStudy (xapproachesinfinity):

when a<0 that is

OpenStudy (xapproachesinfinity):

a being the leading coefficient!